# Strange Stars Admixed with Mirror Dark Matter: Confronting Observations of XTE J1814-338

### v1

@@ 1 INTRODUCTION
Neutron stars (NSs) have long been a subject of intense study due to their extreme physical properties and their role as cosmic laboratories for dense matter physics. Traditionally, NSs are modeled as compact objects composed of nuclear matter, but recent observations have challenged these conventional models. These include measurements of NS masses, radii, and tidal deformabilities from gravitational waves, X-ray timing, and thermonuclear burst oscillations. Notably, NICER observations of PSR J0740+6620 and PSR J0030+0451, along with the exotic source XTE J1814-338, suggest that NSs may have more exotic compositions than previously thought. These observations call for new theoretical models that go beyond standard nuclear matter and explore the possibility of exotic matter inside NSs.
@@
One such possibility is that NSs could actually be strange stars (SSs) (Alcock et al., 1986; Bombaci, 1997; Alcock & Olinto, 1988; Madsen, 1999; Weber, 2005; Zhang et al., 2024), which are made of strange quark matter (SQM). SQM, which consists of up ($u$), down ($d$), and strange ($s$) quarks, was proposed by Witten (1984); Farhi & Jaffe (1984), and others as a potential ground state of matter, even more stable than nuclear matter. SSs could explain the existence of NSs with unusually high masses and small radii, such as PSR J0740+6620, without invoking exotic particles beyond the Standard Model. In this scenario, the entire star is composed of deconfined quarks, unlike conventional NSs which are primarily made of neutrons.
@@
However, SSs alone may not fully account for recent observations. Mirror dark matter (MDM) has emerged as a promising extension to the SS model. MDM, first proposed by Foot et al. (1991) and later expanded by Berezhiani et al. (2021), is a stable, self-interacting dark matter candidate that interacts with ordinary matter only through gravity. The concept of a dark matter halo surrounding an SS could help to explain the peculiar mass and radius observations of XTE J1814-338. This star exhibits an unusually small radius ($R\approx 7$ km) and relatively low mass ($M\approx 1.2M_{\odot}$), which are difficult to reconcile with standard NS or even SS models. We propose that this compact object could be a mirror strange star (MSS), with an SQM core surrounded by an MDM halo. This dual-component model may explain both the high compactness of XTE J1814-338 and the more typical mass-radius relationships observed in other NSs and SSs.
@@
Liu et al. (2024b) explored the impact of dark matter halos on the pulse profiles of X-ray pulsars, showing that dark matter can significantly modify pulse shapes and peak fluxes. This complements our findings by illustrating how dark matter, whether mirror or bosonic, could influence the observable properties of compact objects like SSs. The presence of dark matter halos or admixtures, as demonstrated in their study, could provide additional observational evidence for dark matter in astrophysical systems.
@@
Recently, Pitz & Schaffner-Bielich (2024) proposed that XTE J1814-338 might be a boson star with a nuclear matter core, a model that shares some similarities with the MSS model. However, the boson star model requires the introduction of new scalar particles beyond the Standard Model, which complicates its physical interpretation. Our approach, by contrast, sticks to known particles and symmetries, making it a more conservative extension of the SS hypothesis.
@@
Recent gravitational wave observations from events like GW170817 and GW190814 have provided new constraints on the equation of state (EOS) of NSs. These observations place upper limits on the tidal deformability and radius of NSs, favoring stiff EOSs that are more consistent with SQM or dark matter-admixed stars (Abbott et al., 2018; Dittmann et al., 2024). MDM admixed SSs offer a way to reconcile these constraints with the observed compactness of sources like XTE J1814-338.
@@
In this paper, we aim to build on these ideas by exploring the mass-radius relationships of SSs admixed with MDM, comparing them with recent observations. In particular, we focus on explaining the properties of XTE J1814-338 and other peculiar NS candidates that defy explanation through conventional models. Our study expands upon previous work on SSs (Weber, 2005; Kuerban et al., 2020; Geng et al., 2021) and MDM (Berezhiani et al., 2021; Yang et al., 2021a), providing a unified framework for understanding compact objects with extreme densities and exotic compositions.
@@
This paper is structured as follows. In Sec. 2, we present the theoretical framework for SSs admixed with MDM, focusing on the EOS. Section 3 details the results, analyzing mass-radius relations and parameter constraints. In Sec. 4, we discuss observational implications and propose future tests for the model using gravitational waves and X-ray timing. A summary is provided in Sec. 5.
2 Equation of State of Strange Quark Matter and Mirror Dark Matter
The equation of state (EOS) plays a critical role in determining the internal structure and observable properties of compact stars, such as their mass-radius relationship, stability, and possible phase transitions. In this study, we focus on SQM and MDM, exploring the potential existence of compact stars composed of these exotic forms of matter.
2.1 Strange Quark Matter
@@
SQM is hypothesized to consist of roughly equal numbers of up ($u$), down ($d$), and strange ($s$) quarks, along with a small admixture of electrons to maintain charge neutrality. This composition is predicted by the strange matter hypothesis, originally proposed by Itoh (1970) and further developed in seminal works such as Bodmer (1971), Witten (1984), and Farhi & Jaffe (1984). According to this hypothesis, SQM may be more stable than ordinary nuclear matter, implying that compact stars could exist as SSs rather than NSs.
@@
For the EOS of SQM, we employ the modified bag model, which has been extensively used in studies of SSs (Farhi & Jaffe, 1984; Alcock et al., 1986; Haensel et al., 1986; Weber, 2005). In this model, the quarks are considered to be massless (for $u$ and $d$ quarks), while the $s$ quark has a finite mass ($m_{s}=93$ MeV, as adopted from Navas et al. (2024)). First-order perturbative corrections to the strong interaction coupling constant $\alpha_{S}$ are included to account for interactions among quarks. This model also incorporates the bag constant ($B$), which represents the vacuum pressure that confines quarks inside hadrons or within the SSs.
@@
where $i=u,d,s,e$ and $\Omega_{i}$ are the thermodynamic potentials for up, down, strange quarks, and electrons (the formalism of $\Omega_{i}$ can be found in Alcock et al. (1986) and Yang et al. (2020)), $\mu_{i}$ are the respective chemical potentials. Chemical equilibrium is maintained via weak interaction processes, resulting in the relations:
@@
The modified bag model, while effective in describing SQM, has certain limitations, such as the omission of higher-order quantum corrections and possible phase transitions at extreme densities. However, it remains one of the most widely used models due to its simplicity and ability to reproduce key features of compact stars in high-density regimes (Weber, 2005; Alford et al., 2005; Weissenborn et al., 2011; Bhattacharyya et al., 2016; Zhou et al., 2018; Miao et al., 2021; Zhang & Mann, 2021).
2.2 Mirror Dark Matter
@@
MDM is a candidate for the stable, self-interacting dark matter that could coexist with ordinary matter. This concept arises from parity-symmetric extensions of the Standard Model, first proposed by Lee & Yang (1956) and Kobzarev et al. (1966), and later expanded by Foot et al. (1991). In these models, every Standard Model particle has a mirror counterpart, with the key distinction that mirror particles interact via right-handed interactions, while ordinary particles interact via left-handed interactions. These mirror particles form a hidden, parallel universe, with MDM potentially interacting only gravitationally with ordinary matter.
@@
The theoretical framework of MDM is reviewed in detail by Foot (2004, 2014); Berezhiani (2004, 2018), among others. In this study, we assume the simplest case in which the microphysics of MDM mirrors that of ordinary matter, implying that MDM follows the same EOS as SQM. This means that the mirror SQM (mirror up ($u^{\prime}$), down ($d^{\prime}$), and strange ($s^{\prime}$) quarks) is governed by the same thermodynamic principles as ordinary SQM.
@@
In the minimal parity-symmetric extension of the Standard Model, the gauge group is doubled, $G\otimes G$, where $G$ represents the Standard Model gauge group. This symmetry ensures that the interactions within the mirror sector are identical to those in the ordinary sector, except for the handedness of the interactions. Consequently, the EOS for MDM is identical to that of SQM.
2.3 Interaction Between SQM and MDM
@@
In our model, SQM and MDM interact solely through gravitational forces. No direct interaction between quarks and mirror quarks is assumed, as the nature of such interactions remains speculative and unstudied. However, if quark-mirror quark interactions exist, they are expected to be weak enough to be neglected in our study, as suggested by Berezhiani et al. (2021). Therefore, we model the star as a two-component system, where the gravitational coupling between the SQM core and the MDM halo governs the overall structure and dynamics of the star.
@@
This two-fluid formalism has been widely employed in previous studies of compact stars with dark matter components (e.g., Sandin & Ciarcelluti, 2009; Ciarcelluti & Sandin, 2011; Leung et al., 2011; Li et al., 2012; Mukhopadhyay & Schaffner-Bielich, 2016; Panotopoulos & Lopes, 2017, 2018; Ellis et al., 2018a, b; Nelson et al., 2019; Garani et al., 2019; Ivanytskyi et al., 2020; Lee et al., 2021; Kain, 2021; Ciancarella et al., 2021; Yang et al., 2021a; Karkevandi et al., 2022; Miao et al., 2022; Emma et al., 2022; Diedrichs et al., 2023; Ferreira & Fraga, 2023; Hippert et al., 2023; Giangrandi et al., 2023; Rezaei, 2023; Liu et al., 2023; Shakeri & Karkevandi, 2024; Hong & Ren, 2024; Karkevandi et al., 2024; Sun & Wen, 2024; Liu et al., 2024a; Thakur et al., 2024; Barbat et al., 2024; Zhen et al., 2024; Shawqi & Morsink, 2024; Liu et al., 2024b). In these studies, the dark matter component is treated as a separate fluid interacting gravitationally with the ordinary matter. The same approach is used in our model, which treats the SQM and MDM components as independent fluids, each obeying its own EOS but interacting through their mutual gravitational field. This approach allows us to explore the impact of MDM on the mass-radius relation of compact stars, particularly in the case of XTE J1814-338, which exhibits unusual mass and radius measurements.
@@ 3 Results and Discussion
For the given EOS of SQM, we compute the structure of SSs admixed with MDM using the two-fluid formalism (Yang et al., 2021a). This formalism, widely employed in the study of compact stars containing dark matter, assumes that the SQM and MDM components interact only through gravity, with no direct interactions between quarks and mirror quarks. Such a framework allows us to isolate the gravitational influence of the dark matter component and study its impact on the mass-radius relation of SSs.
3.1 Mass-Radius Relation of SSs without MDM
@@
In Fig. 1, we present the mass-radius relation of pure SSs (i.e., without MDM) for different values of the bag constant $B^{1/4}$ and a strong interaction coupling constant of $\alpha_{S}=0.6$. We define the mass fraction of MDM as $f_{D}\equiv M_{D}/M$, where $M$ is the total mass of the star and $M_{D}$ is the mass of the MDM component. Fig. 1 corresponds to $f_{D}=0$, meaning no MDM is included.
@@
Figure 1: The mass-radius relation of SSs without MDM (i.e., with a mass fraction of MDM $f_{D}=0$) for $\alpha_{S}=0.6$ and different values of $B^{1/4}$ is shown. The pink and cyan regions represent the mass and radius of PSR J0740+6620 as presented by Cromartie et al. (2020); Fonseca et al. (2021); Dittmann et al. (2024) and Salmi et al. (2024), respectively. The blue and green regions show the mass and radius of PSR J0030+0451 (Vinciguerra et al., 2024) and PSR J0437-4715 (Choudhury et al., 2024), respectively. The grey region indicates the mass and radius estimates for the central compact object within the supernova remnant HESS J1731-347 (Doroshenko et al., 2022). The red region shows data from the observation of XTE J1814-338 (Kini et al., 2024).
@@
The range of values for $B^{1/4}$ between 128.6 MeV and 141.9 MeV was chosen because they satisfy both the "2-flavor line" and "3-flavor line" constraints, as shown in Fig. 2. These constraints ensure that the SQM remains stable and energetically favorable in comparison to nuclear matter, which is essential for the existence of SSs.
@@
Figure 2: Constraints on $B^{1/4}$ and $\alpha_{S}$ for SQM. The solid, dashed, and dotted grey lines correspond to $M_{\rm max}=2.0\,M_{\odot},2.2\,M_{\odot},$ and $2.4\,M_{\odot}$, respectively. The green line represents $\Lambda(1.4)=580$. All the curves for mass, radius, and tidal deformability are calculated for pure SSs without MDM, i.e., for $f_{D}=0$.
@@
From Fig. 1, we observe that the green, red, and blue curves match most of the current observational data. However, they fail to explain the exotic mass and radius observation of XTE J1814-338. Specifically, the green line marginally satisfies the mass and radius constraints for PSR J0437-4715 (Choudhury et al., 2024), while the blue line only marginally satisfies the observation of PSR J0740+6620 (Salmi et al., 2024). For $\alpha_{S}=0.6$, the results suggest that the allowed values of $B^{1/4}$ must fall within the range of 131.6 MeV to 140.7 MeV to satisfy the majority of observations.
3.2 Parameter Space and Constraints

### v4

@@ I INTRODUCTION
Neutron stars (NSs) have long been a subject of intense study due to their extreme physical properties and their role as cosmic laboratories for dense matter physics. Traditionally, NSs are modeled as compact objects composed of nuclear matter. However, recent observations have challenged these conventional models, including measurements of NS masses, radii, and tidal deformabilities derived from gravitational waves, X-ray timing, and thermonuclear burst oscillations. Notably, NICER observations of PSR J0740+6620 and PSR J0030+0451, along with the exotic source XTE J1814-338, suggest that NSs may possess more exotic compositions than previously thought. In particular, the compact stellar object XTE J1814-338 defies conventional explanations. This object exhibits an unusually low radius (7($\sim$ 7( ∼ 7 km) and mass ($\sim 1.2\,M_{\odot}$). These observations call for new theoretical models that extend beyond standard nuclear matter and explore the possibility of exotic matter within NSs.
@@
One such possibility is that NSs could actually be strange stars (SSs) [1, 2, 3, 4, 5, 6], which are made of strange quark matter (SQM). SQM, which consists of up ($u$), down ($d$), and strange ($s$) quarks, was proposed by Witten [7], Farhi and Jaffe [8], and others as a potential ground state of matter, even more stable than nuclear matter. SSs could explain the existence of NSs with unusually high masses and small radii, such as PSR J0740+6620, without invoking exotic particles beyond the Standard Model. In this scenario, the entire star is composed of deconfined quarks, unlike conventional NSs which are primarily made of neutrons.
@@
However, SSs alone may not fully account for recent observations. Mirror dark matter (MDM) has emerged as a promising extension to the SS model. MDM, first proposed by Foot et al. [9] and later expanded by Berezhiani et al. [10], is a stable, self-interacting dark matter candidate that interacts with ordinary matter only through gravity. The concept of a dark matter halo surrounding an SS could help to explain the peculiar mass and radius observations of XTE J1814-338 [11]. This star exhibits an unusually small radius ($R\approx 7$ km) and relatively low mass ($M\approx 1.2M_{\odot}$), which are difficult to reconcile with standard NS or even SS models. We propose that this compact object could be a mirror strange star (MSS), with an SQM core surrounded by an MDM halo. This dual-component model may explain both the high compactness of XTE J1814-338 and the more typical mass-radius relationships observed in other NSs and SSs.
@@
Liu et al. [12] explored the impact of dark matter halos on the pulse profiles of X-ray pulsars, showing that dark matter can significantly modify pulse shapes and peak fluxes. This complements our findings by illustrating how dark matter, whether mirror or bosonic, could influence the observable properties of compact objects like SSs. The presence of dark matter halos or admixtures, as demonstrated in their study, could provide additional observational evidence for dark matter in astrophysical systems.
@@
Recently, Pitz and Schaffner-Bielich [13] proposed that XTE J1814-338 might be a boson star with a nuclear matter core, a model that shares some similarities with the MSS model. However, the boson star model requires the introduction of new scalar particles beyond the Standard Model, which complicates its physical interpretation. Our approach, by contrast, sticks to known particles and symmetries, making it a more conservative extension of the SS hypothesis.
@@
Recent gravitational wave observations from events like GW170817 and GW190814 have provided new constraints on the equation of state (EOS) of NSs. These observations place upper limits on the tidal deformability and radius of NSs, favoring stiff EOSs that are more consistent with SQM or dark matter-admixed stars [14, 15, 16]. MDM admixed SSs offer a way to reconcile these constraints with the observed compactness of sources like XTE J1814-338.
@@
In this paper, we aim to build on these ideas by exploring the mass-radius relationships of SSs admixed with MDM, comparing them with recent observations. In particular, we focus on explaining the properties of XTE J1814-338 and other peculiar NS candidates that defy explanation through conventional models. Our study expands upon previous work on SSs [5, 17, 18] and MDM [10, 19], providing a unified framework for understanding compact objects with extreme densities and exotic compositions.
@@
This paper is structured as follows. In Sec. II, we present the theoretical framework for SSs admixed with MDM, focusing on the EOS. Section III details the results, analyzing mass-radius relations and parameter constraints. In Sec. IV, we discuss observational implications and propose future tests for the model using gravitational waves and X-ray timing. A summary is provided in Sec. V.
II Equation of State of Strange Quark Matter and Mirror Dark Matter
The equation of state (EOS) determines the relationship between pressure and energy density, serving as the foundation for modeling the properties of compact stars, such as their mass-radius relationship, stability, and possible phase transitions. In this study, we employ a modified bag model for SQM, which accounts for key properties such as charge neutrality and chemical equilibrium. To incorporate MDM, we assume a symmetric EOS for mirror quark matter, reflecting the same thermodynamic principles as ordinary SQM. This unified EOS framework allows us to analyze the effects of MDM on the mass-radius relationship and test its alignment with observational constraints.
II.1 Strange Quark Matter
@@
SQM is hypothesized to consist of roughly equal numbers of up ($u$), down ($d$), and strange ($s$) quarks, along with a small admixture of electrons to maintain charge neutrality. This composition is predicted by the strange matter hypothesis, originally proposed by Itoh [20] and further developed in seminal works such as [21, 7, 8]. According to this hypothesis, SQM may be more stable than ordinary nuclear matter, implying that compact stars could exist as SSs rather than NSs.
@@
For the EOS of SQM, we employ the modified bag model, which has been extensively used in studies of SSs [8, 1, 22, 5]. In this model, the quarks are considered to be massless (for $u$ and $d$ quarks), while the $s$ quark has a finite mass ($m_{s}=93$ MeV [23]). First-order perturbative corrections to the strong interaction coupling constant $\alpha_{S}$ are included to account for interactions among quarks. This model also incorporates the bag constant ($B$), which represents the vacuum pressure that confines quarks inside hadrons or within the SSs.
@@
where $i=u,d,s,e$; $\Omega_{i}$ are the thermodynamic potentials for up, down, strange quarks, and electrons (the formalism of $\Omega_{i}$ can be found in Refs. [1, 24]), and $\mu_{i}$ are the respective chemical potentials. Chemical equilibrium is maintained via weak interaction processes, resulting in the relations:
@@
The modified bag model, while effective in describing SQM, has certain limitations, such as the omission of higher-order quantum corrections and possible phase transitions at extreme densities. However, it remains one of the most widely used models due to its simplicity and ability to reproduce key features of compact stars in high-density regimes [5, 25, 26, 27, 28, 29, 30, 31, 32].
II.2 Mirror Dark Matter
@@
MDM is a candidate for the stable, self-interacting dark matter that could coexist with ordinary matter. This concept arises from parity-symmetric extensions of the Standard Model, first proposed by Lee and Yang [33] and Kobzarev et al. [34], and later expanded by others [35, 36, 37, 9, 38, 39]. In these models, every Standard Model particle has a mirror counterpart, with the key distinction that mirror particles interact via right-handed interactions, while ordinary particles interact via left-handed interactions. These mirror particles form a hidden, parallel universe, with MDM potentially interacting only gravitationally with ordinary matter.
@@
The theoretical framework of MDM is reviewed in detail in Refs. [40, 41, 42, 43], among others. In this study, we assume the simplest case in which the microphysics of MDM mirrors that of ordinary matter, implying that MDM follows the same EOS as SQM. This means that the mirror SQM (mirror up ($u^{\prime}$), down ($d^{\prime}$), and strange ($s^{\prime}$) quarks) is governed by the same thermodynamic principles as ordinary SQM.
@@
In the minimal parity-symmetric extension of the Standard Model, the gauge group is doubled, $G\otimes G$, where $G$ represents the Standard Model gauge group. This symmetry ensures that the interactions within the mirror sector are identical to those in the ordinary sector, except for the handedness of the interactions. Consequently, the EOS for the mirror SQM is identical to that of SQM.
Recent studies, such as those by Liu et al. [44] and Mariani et al. [45], have demonstrated the critical role of the dark matter self-interaction cross-section constraint ($\sigma/m$) in determining the observable properties of compact stars. This constraint, derived from galaxy cluster observations, provides a range ($\sigma/m\sim 0.1-10\,\mathrm{cm^{2}/g}$) [46] that aligns with both astrophysical data and theoretical models.
For MDM, the cross-section constraint derived from galaxy clusters observations might be related to the QCD scale of it ($\Lambda^{\prime}$). Mohapatra et al. [39] found that for scenarios with appreciable mirror symmetry breaking, e.g., the QCD scale of the MDM ($\Lambda^{\prime}$) is 30 times the ordinary matter ($\Lambda$), the mirror hydrogen atoms could satisfy the cross-section constraint derived from galaxy clusters observations. In this case, the maximum mass and the radius of the pure MSSs will be smaller than ordinary SSs (just like the case of mirror NSs [47]), and the structure of the MDM admixed SSs will also be changed significantly (like MDM admixed NSs [48]).
However, Mohapatra and Nussinov [49] found that to fulfill the observation of $n-n^{\prime}$($n^{\prime}$ is the mirror neutron) oscillations, a very precise mirror symmetry is required (same result is found by Berezhiani [50]), which leads the result that the cross-section of mirror hydrogen atoms is much higher than the upper bound from galaxy clusters observations. Nevertheless, Mohapatra and Nussinov [49] proposed that if most of the MDM in galaxies and galaxy clusters reside in collisonless stars and not in gas, the cross-section constraint could be satisfied for mirror scenarios with precise mirror symmetry.
Thus, considering the current status about the study of the cross-section constraint to MDM, we will use the symmetric model ($\Lambda^{\prime}=\Lambda$, where the EOS for the mirror SQM is the same as that of ordinary SQM) following Mohapatra and Nussinov [49] and Berezhiani [50] in this paper.
II.3 Interaction between SQM and MDM
@@
In our model, SQM and MDM interact solely through gravitational forces. Although the neutron–mirror neutron mixing has been widely studied [51, 52, 53, 54, 55], the nature of direct interaction between quarks and mirror quarks remains speculative and unstudied. However, if quark-mirror quark interactions exist, they are expected to be weak enough to be neglected in our study, as suggested by Berezhiani et al. [10]. Therefore, we model the star as a two-component system, where the gravitational coupling between the SQM core and the MDM halo governs the overall structure and dynamics of the star.
@@
This two-fluid formalism has been widely employed in previous studies of compact stars with dark matter components [e.g., 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 48, 80, 81, 82, 83, 84, 45, 85, 86, 87, 44, 88, 89, 90, 91, 92, 93]. In these studies, the dark matter component is treated as a separate fluid interacting gravitationally with the ordinary matter. The same approach is used in our model, which treats the SQM and MDM components as independent fluids, each obeying its own EOS but interacting through their mutual gravitational field. This approach allows us to explore the impact of MDM on the mass-radius relation of compact stars, particularly in the case of XTE J1814-338, which exhibits unusual mass and radius measurements.
@@ III Results and Discussion
For the given EOS of SQM, we compute the structure of SSs admixed with MDM using the two-fluid formalism [19]. This formalism, widely employed in the study of compact stars containing dark matter, assumes that the SQM and MDM components interact only through gravity, with no direct interactions between quarks and mirror quarks. Such a framework allows us to isolate the gravitational influence of the dark matter component and study its impact on the mass-radius relation of SSs.
III.1 Mass-Radius Relation of SSs without MDM
@@
In Fig. 1, we present the mass-radius relation of pure SSs (i.e., without MDM) for different values of the bag constant $B^{1/4}$ and a strong interaction coupling constant of $\alpha_{S}=0.6$. We define the mass fraction of MDM as $f_{D}\equiv M_{D}/M$, where $M$ is the total mass of the star and $M_{D}$ is the mass of the MDM component. Figure 1 corresponds to $f_{D}=0$, meaning no MDM is included.
@@
Figure 1: The mass-radius relation of SSs without MDM (i.e., with a mass fraction of MDM $f_{D}=0$) for $\alpha_{S}=0.6$ and different values of $B^{1/4}$ is shown. The pink and cyan regions represent the mass and radius of PSR J0740+6620 as presented in Refs. [94, 95, 16] and Ref. [96], respectively. The blue and green regions show the mass and radius of PSR J0030+0451 [97] and PSR J0437-4715 [98], respectively. The grey region indicates the mass and radius estimates for the central compact object within the supernova remnant HESS J1731-347 [99]. The red region shows data from the observation of XTE J1814-338 [11]. For comparison, two lines for NSs are also presented: the dashed line corresponds to APR3 EOS [100], and the dash-dotted line corresponds to KWCZ3 EOS (for the case of $C_{\sigma}=14$, $L=40$) [101].
@@
The range of values for $B^{1/4}$ between 128.6 MeV and 141.9 MeV was chosen because they satisfy both the ”2-flavor line” and ”3-flavor line” constraints, as shown in Fig. 2. These constraints ensure that the SQM remains stable and energetically favorable in comparison to nuclear matter, which is essential for the existence of SSs.
@@
Figure 2: Constraints on $B^{1/4}$ and $\alpha_{S}$ for SQM. The solid, dashed, and dotted grey lines correspond to $M_{\rm max}=2.0\,M_{\odot},2.2\,M_{\odot},$ and $2.4\,M_{\odot}$, respectively. The green line represents $\Lambda(1.4)=580$ [$\Lambda(1.4)$ is the dimensionless tidal deformability of a 1.4 $M_{\odot}$ pure SS]. All the curves for mass, radius, and tidal deformability are calculated for pure SSs without MDM, i.e., for $f_{D}=0$.
@@
From Fig. 1, we observe that the green, red, and blue curves match most of the current observational data. However, they fail to explain the exotic mass and radius observation of XTE J1814-338. Specifically, the green line marginally satisfies the mass and radius constraints for PSR J0437-4715 [98], while the blue line only marginally satisfies the observation of PSR J0740+6620 [96]. For $\alpha_{S}=0.6$, the results suggest that the allowed values of $B^{1/4}$ must fall within the range of 131.6 MeV to 140.7 MeV to satisfy the majority of observations.
III.2 Parameter Space and Constraints